10 grams of solid iodine was mixed with 8 centimetre cube of ethanol and 2.5 grams remained and dissolved after thorough stirring for 5 minutes determine the rate of reaction in kilogram per decimeter cube
1
Expert's answer
2020-03-27T09:01:32-0400

1) "\Delta m = 10-2.5 = 7.5 g"

"7.5 g (\frac{1 kg}{1000 g}) = 7.5 \times 10^{-3} kg"


2) "8 cm^3 (\frac{1 dm^3}{1000 cm^3} )=8\times 10^{-3} dm^3"


3) "Rate = \frac {\Delta m}{V \times \Delta t} = \frac{7.5\times10^{-3} kg}{8\times10^{-3} dm^3\times 5 min}= 0.19 \frac{kg}{dm^3\times min}"